2. Two fair six-sided dice are tossed independently. Let M = the maximum of the two tosses (so M (1,5) = 5, M (3,3) = 3, etc). (a) I4 ptsl Find the probability mass function of M. (b) 14 pts] Find the cumulative distribution function of M and graph it. (c) 12 pts] Find the expected value of M (d) 12 pts] Find the variance of M. (e) 12 pts] Find the standard deviation of M.
Answer #1

Solution:

   we are given that : two fair six sided dice are tossed independently.

M = Maximum of the two tosses.

Thus we get:

First Die Second Die M = Maximum of two dice
1 1 1
1 2 2
1 3 3
1 4 4
1 5 5
1 6 6
2 1 2
2 2 2
2 3 3
2 4 4
2 5 5
2 6 6
3 1 3
3 2 3
3 3 3
3 4 4
3 5 5
3 6 6
4 1 4
4 2 4
4 3 4
4 4 4
4 5 5
4 6 6
5 1 5
5 2 5
5 3 5
5 4 5
5 5 5
5 6 6
6 1 6
6 2 6
6 3 6
6 4 6
6 5 6
6 6 6

Part a) Probability mass function of M.

We need to find frequencies of each possible values of M and then divide each frequency by N = 36

M f = frequency P(M)
1 1 0.0278
2 3 0.0833
3 5 0.1389
4 7 0.1944
5 9 0.2500
6 11 0.3056
N = 36

Part b) Cumulative distribution function of M and Graph it.

To get cumulative distribution function , we need to find F(M) column

That is find cumulative sum of probabilities of each value of M.

Thus we get :

M P(M) Calculations of F(M) F(M)
1 0.0278                        0.0278 0.0278
2 0.0833 0.0278+0.0833=0.1111 0.1111
3 0.1389 0.1111+0.1389=0.2500 0.2500
4 0.1944 0.2500+0.1944=0.4444 0.4444
5 0.2500 0.4444+0.2500=0.6944 0.6944
6 0.3056 0.6944+0.3056=1.0000 1.0000

CDF of F(M) 1.2 1.0000 0.8 0.6944 2 0.6 0.4444 0.4 0.2500 0.0278 4

Part c) Expected value of M

E(M)= \sum M \times P(M)

M P(M) M * P(M)
1 0.0278 0.0278
2 0.0833 0.1667
3 0.1389 0.4167
4 0.1944 0.7778
5 0.2500 1.2500
6 0.3056 1.8333
\sum M \times P(M)=4.4722

Thus

E(M)= \sum M \times P(M)

E(M)=4.4722

Part d) Variance of M

Var(M)=E(M^{2})-(E(M))^{2}

Where

E(M^{2})=\sum M^{2} \times P(M)

M P(M) M * P(M) M^2 * P(M)
1 0.0278 0.0278 0.0278
2 0.0833 0.1667 0.3333
3 0.1389 0.4167 1.2500
4 0.1944 0.7778 3.1111
5 0.2500 1.2500 6.2500
6 0.3056 1.8333 11.0000
\sum M \times P(M)=4.4722 \sum M^{2} \times P(M)=21.9722

Thus E(M^{2})=\sum M^{2} \times P(M)

E(M^{2})=21.9722

Var(M)=E(M^{2})-(E(M))^{2}

Var(M)=21.9722 -(4.4722 )^{2}

Var(M)=21.9722 -20.0008

Var(M)=1.9715

Part e) Standard deviation of M

SD(M)=\sqrt{Var(M)}

SD(M)=\sqrt{1.9715 }

SD(M)=1.4041